OpenStudy (anonymous):

Special Cases of Factoring, need the steps please. 1. y2-49 2. 4x2-25 3. 81x2 – 1 4. 16y2-8y+1

OpenStudy (anonymous):

y^2-49= (y^2-7^2)=(y+7)(y-7) 4x^2-25=((2x)^2-5^2)=(2x+5)(2x-5) .................

OpenStudy (anonymous):

remember the formula: \[a ^{2}-b ^{2}=(a-b)(a+b)\]

OpenStudy (anonymous):

use it